HUBBLE SPACE TELESCOPE DAILY REPORT #1681 PERIOD COVERED: 0000Z (UTC) 07/23/96 - 0000Z (UTC) 07/24/96 Daily Status Report as of 206/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/RD 6595 (Does Pluto Have a Haze Layer?) The Faint Object Spectrograph (Red Side) was used to make = spectrographic observations of the planet Pluto in order to determine if = the planet has a haze layer. These observations were made during a = series of stellar occultations. The objective is to study the physical = structure of the atmosphere at kilometer-level spatial resolution. = During this proposal there occurred a status buffer message at = 205/024015Z that described an FOS target aquisition failure which had = been scheduled at 205/023522Z. Aperture mapping had not found a target. = This happened about 30 minutes into a three-hour observation. HSTAR 5661 = was written. Otherwise, the observation completed as scheduled, and no = further anomalies were noted. 1.2 Completed WF/PC-2 6187 (WF/PC-2 Cycle 5 Calibration: Earth Flat) The WF/PC-2 was used to make Earth streak flat calibration images. = These streak flats will be used to produce flat fields, to map the OTA = illumination pattern and, in combination with other data, to produce new = superflats. The observations were completed as planned, and no problems = were reported. 1.3 Completed WF/PC-2 6431 (Determination of the Extragalactric = Distance Scale) The WF/PC-2 was used to make observations of the galaxy NGC-4535 = to be used in measuring the periods of Cepheid variable stars within this = galaxy. This will allow for a more accurate estimate of the distances to = these galaxies through the period-luminosity relation for the Cepheid = stars. The observations were executed as scheduled, and no anomalies = were noted. 1.4 Completed Two Sets of FOS/BL 6166 (FOS Cycle 5: Location of = Spectra and Wavelength Calibration) The Faint Object Spectrograph (Blue Side) was used to make = wavelength calibration images to measure the stability of the FOS Blue = Side wavelength scale and spectral locations. A Taled calibration = observations was also taken. The observations were completed as planned, = and no anomalies were reported. 1.5 Completed Two Sets of GHRS/2 6213 (GHRS Aperture Offsets- Cycle 5) The High Resolution Spectrograph (Side Two) was used to make = calibration images of the star HD155763 in order to determine the offsets = of the instruments focal plain apertures. The purpose of this calibration= program is to measure the relative wavelength offsets between the Large = Science Aperture and Small Science Aperture with various gratings. A = wavelength calibration image was also taken. The observations were = completed as scheduled, and no anomalies were reported. 1.6 Completed FOS/RD 6204 (FOS Cycle 5 Internal/External Offset) The Faint Object Spectrograph (Red Side) was used to make = spectrographic calibration images of the object NGC-6833 in order to = measure the spectrograph's internal and external offsets. The internal = lamp was intermingled with the external observations. The observations = were completed as planned, and no problems were reported. 1.7 Completed WF/PC-2 6359 (Core Properties of the Bulges of Spiral = Galaxies) The WF/PC-2 was used to make snapshot survey observations of = spiral galaxy NGC-5879 in order to study its central bulge. The aim is = to investigate the core properties, correlation's with global galactic = properties, and the frequency of occurrence of multiple and/or = miscentered nuclei. The observations were completed as scheduled, and no = anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 2 Successful : 2 2.3 Operations Notes: Operations personnel have ensured that all STR playbacks have been = collected using manually scheduled TDRSS services. In preparation for the simulation being conducted today, HST = operations transferred to the SAMS back-up facility at 206/0245Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: GSFC Internal Simulation #8 (EVAs 3/4) will be conducted today = from 8:00 a.m. until approximately 8:00 p.m. /DMH